Is {sinx, cosx, sin(x+π/4)} a linearly independent set over R ? Justify our answer


1
Expert's answer
2021-02-26T04:57:50-0500

Yes the set "( sinx,cosx,sin(x+\dfrac{\pi}{4}))" is linearly independent over R

Because,


For the linear independency,


"(asinx+bcos+csin(x+\dfrac{\pi}{4})) = 0" for only "a= b=c=0"


Hence, we get only one value of "a, b" and "c" for which it is zero i.e. 0


Therefore, the set is linearly independent.